3. x(t) - y(t) = -2t x(t) + y(t) = 0 x(O) = 0; y(0)=0; x0) = 0; y(0) = 0

Please show all work.
Answer #1

(3) x H) - 1 (4) 1 - 2 xlt) + y 14) = 0 - () . _ (ii) x 10) = 0, 710)=0, x(0)=0, 7(0) = 0. taking laplace toansform of equfrom equation (in) we get -537/5) - Sy(s) = - 2 / 2 5 (52+1) 715) = Cs) = 31541) By the partial fraction methods A52/52+1) +Hence, 769) = 2 + + toans foron both sides laplace taking inverse we we get a 4 Lt) = -2i+t2 + acost ylt) = 2t - 2 sint y(tTherefore, xit) = -2t + sint ylt) = -2+ + 2 + 2 cost